### v1

# Deep learning reveals key predictors of thermal conductivity in covalent organic frameworks

Abstract
The thermal conductivity of covalent organic frameworks (COFs), an emerging class of nanoporous polymeric materials, is crucial for many applications, yet the link between their structure and thermal properties is not well understood. From a dataset of over 2,400 COFs, we find that conventional features like density, pore size, void fraction, and surface area do not reliably predict thermal conductivity. To overcome this, we train an attention-based machine learning model that accurately predicts thermal conductivities, even for structures outside the training set. We then use the attention mechanism to understand why the model works. Surprisingly, dangling molecular branches emerge as key predictors of thermal conductivity, alongside conventional geometric descriptors like density and pore size. Our findings show that COFs with dangling functional groups exhibit lower thermal transfer capabilities than otherwise. Molecular dynamics simulations confirm this, revealing significant mismatches in the vibrational density of states due to the presence of dangling branches.
keywords:
Keywords
Introduction
Figure 1: Overview of this study. A COF structure consists of molecular substructures called linkers and knots arranged in a periodic pattern described by the topology. Through molecular dynamics, we construct a large dataset of 2,471 COFs (with diverse linkers, knots, and topologies) and their corresponding thermal conductivities. We demonstrate that conventional descriptors, such as density, pore size, void fraction, and surface area, fail to predict the thermal conductivity reliably. We employ a machine learning framework based on attention mechanism and transformer architecture to uncover a novel predictor, thus enhancing our understanding of the structure-property relationship of thermal conductivity of COFs.
Covalent organic frameworks[1, 2] (COFs) are an emerging class of nanoporous polymeric materials. Compared to metal-organic framework (MOFs)[3, 4] and zeolites[5], the crystalline backbone of COFs is composed of organic building blocks – known as knots and linkers – connected by strong covalent bonds, thereby offering higher stability[6]. The crystalline nature of COFs, along with their high surface areas, tunable pore sizes, and functionalizable organic linkers, make them exceptionally suited for a wide range of promising applications. These applications include (but are not limited to) photoconductivity,[7, 8] chemo-sensing,[9, 10] catalysis,[11] drug delivery,[12, 13] thermoelectrics,[14, 15] semiconductors,[16] gas storage and separation.[17, 18]
A key property dictating the applications of COFs is thermal transfer. For example, low thermal conductivity is desired for thermoelectrics to maintain large internal thermal gradients and increase efficiency.[19] On the other hand, a high thermal conductivity is desired for gas adsorption and separation where efficient thermal transfer is important for the longevity and stability of the nanoporous membranes.[20] COFs are promising nanoporous candidates for such applications as the ability to design their crystalline topology (e.g., lattice type, pore size) and chemistry via choice of knots and linkers opens up a large and diverse space of thermal conductivities.[21, 22, 23, 1, 2] Therefore, it is of great interest to gain a deeper understanding of the thermal transfer mechanism and to obtain strong structure-property trends, which in turn enables the application-specific design of COFs.
Two approaches can be used to elucidate the thermal structure-property relationships of COFs: experimental trial-and-error methods which involve synthesis and characterization,[24, 25] and computational high-throughput screening which relies on first-principle calculations.[26, 27, 28] With COFs offering practically an unlimited design space, an experimental trial-and-error approach to screen new COF candidates (including developing new synthesis routes for each candidate) is prohibitively inefficient. Moreover, experimentally synthesized COFs will always contain crystalline defects,[29, 30] which greatly influence the thermal conductivity. This makes it difficult to relate the thermal transfer mechanisms to the geometrical and chemical make-up of a COF structure. On the contrary, molecular dynamics (MD) simulations [31] offer a high-throughput virtual screening alternative to lab-based experiments. However, despite recent advances in computing hardware, even virtual screening can be prohibitively inefficient for a large design space. This is highlighted by the example that generating a dataset of just 2,471 two-dimensional COFs and their thermal conductivity for this study required 1.3 million CPU-hours in cloud computing (translating into approximately four months and \$70,000 in time and cost, respectively). This underscores the need for efficient structure-property maps that are accurate, generalizable, and interpretable.
Recently, Islamov et al.[32] performed a high-throughput screening study of over 10,000 MOF structures and their thermal conductivity ($\kappa$) using MD. They found that the majority of the MOFs possess $\kappa<1\ \text{W}\text{m}^{-1}\text{K}^{-1}$ with a few exceptions possessing ultra-high thermal conductivity ($\kappa>10\ \text{W}\text{m}^{-1}\text{K}^{-1}$). While COFs are generally more thermally stable (owing to the strong covalent bonding of the building blocks), a recent study by Thakur et al.[33] on high-throughput screening study on over 10,000 COFs demonstrated that, generally, COFs also possess a similar range of $\kappa$, with the majority of the structures exhibiting $\kappa<1\ \text{W}\text{m}^{-1}\text{K}^{-1}$. Interestingly, MOFs and COFs possess similar structure-property trends. (i) Increasing the pore size is mostly sufficient to achieve a low thermal conductivity. (ii) Additionally, one can increase the void fraction, the surface area, or include heavy atoms to further decrease thermal conductivity. (iii) However, to increase the thermal conductivity, there are many factors that need to align. Islamov et al.[32] demonstrated that while lowering the pore size and increasing the density results in a higher ceiling of attainable thermal conductivity, other factors, such as the topology, mass-mismatch, and linker length also need to be accounted for. Thakur et al.[33] came to similar conclusions, alongside the observation that aligning the polymeric chains in the structure to heat flow direction raises the ceiling of attainable thermal conductivity. However, these hand-crafted guidelines and ad hoc correlations are insufficient for developing an accurate and generalizable predictive model that captures the thermal conductivity structure-property relationships of COFs. Our study shows that no combination of commonly used descriptors consistently predicts thermal conductivity. Consequently, a definitive method for identifying COFs with tailor-made thermal properties for various applications remains elusive.
To address this knowledge gap, we turn to machine learning (ML) with an emphasis on interpretability and explainability. There are primarily two routes for ML modeling of structure-property maps of porous polymers (including but not limited to COFs) –
  • high intepretability, limited accuracy: using hand-crafted and pre-extracted high-level of the crystalline network (e.g., pore size, density, surface area, atomic composition) as input to classical regression algorithms;[34, 35, 36]
  • high accuracy, limited interpretability: using information-rich but raw graph representation of the crystalline network (where nodes denote atoms with features such as position and atom type, and edges denote bonds with the bond order as the feature) as input to end-to-end deep learning-based regression algorithms.[37, 38, 39, 40]
Here, we bridge the two routes to develop an accurate predictive model using deep learning while also combining interpretability insights from deep learning with prior knowledge and descriptors to explain the thermal conductivity structure-property relationships of COFs (see Figure 1 for an overview).
In the following, we first conduct a large-scale data analysis of the thermal conductivity structure-property relations of COFs and identify the deficiencies in commonly used descriptors. Next, we provide a deep learning model that accurately predicts the thermal conductivity of COFs. Subsequently, an interpretability analysis of the deep learning model uncovers the presence of dangling atoms in the crystalline network of COFs as a strong and so-far missing key predictor for thermal conductivity. Additional physics-based analysis sheds light on the important role of dangling atoms in lowering the thermal conductivity of COFs by disrupting heat transfer pathways. We close by utilizing the ML model for efficient high-throughput screening and identifying COFs with extreme thermal conductivities.
Results
High-throughput data analysis
We create a labeled dataset of COFs and their thermal conductivities by selecting 2,471 two-dimensional COFs from the unlabeled dataset of Mercado et al.[41] The COF structures in this design space are made up of one type of nodal linker (also referred to as knot) and one type of connecting linker. The sampled subset consists of COFs with 104 different linkers arranged in 25 different topologies (e.g., honeycomb, kagome, square lattice, etc.), and bonded by four different types of covalent linkages (i.e., carbon-carbon, amine, amide, and imine). We conduct non-equilibrium molecular dynamics (NEMD) simulations to calculate thermal conductivities in two orthogonal in-plane directions (see Supplementary Information Section 1.1 for more details). In Supplementary Information Section 1.2, we show that there is minimal in-plane anisotropy in thermal conductivity across the dataset. Therefore, we utilize average thermal conductivity in both directions, denoted as $\kappa$, as the quantity of interest to explore the structure-property relationships of COFs. In the following and in Figure 2, we present the observations from data analysis of the key descriptors of COFs and thermal conductivities in our labelled dataset.
Figure 2: Distribution of (a) $\kappa$ versus density (with the color indicating data count per bin), (b) $\kappa$ versus largest pore diameter (LPD), (c) $\kappa$ versus void fraction, and (d) $\kappa$ versus gravimetric surface area (GSA). Also shown are four COF structures (i-iv) with contrasting properties, marked by green triangles in the plots above.
Density: Figure 2a illustrates the distribution of $\kappa$ vs. density. There is a large spread of densities within the dataset and the majority of the COFs possess a density of 0.45-0.55 $\mathrm{g\ cm^{-3}}$. Moreover, in accordance with previous findings,[33] we observe an increasing trend in $\kappa$ with increasing density. However, increasing the density increases the ceiling of attainable $\kappa$, but does not guarantee a high value.
Pore size: The pore size of any COF is represented by the largest pore diameter (LPD). Figure 2b shows that LPD in the dataset is spread over a large range between 5 Å and 80 Å and a concentration of structures in the 15-25 Å region. Moreover, there is an inverse relationship between the pore size and the thermal conductivity, as previously discovered by Freitas et al.,[42] where the range of achievable $\kappa$ decreases significantly with increasing pore sizes. Here again, a small pore size does not guarantee a high thermal conductivity. On the other hand, the pore size can be taken as the sole design parameter to effectively reduce the thermal conductivity range (as mentioned in refs.[33, 32]).
Void fraction: Figure 2c shows that the distribution of the void fractions of COFs across the dataset are within the 0.5–0.96 range, with the biggest share of COFs possessing a void fraction of around 0.85 and higher. Furthermore, a high void fraction of around 0.85 delivers both the largest values and range of attainable $\kappa$ within our dataset. This finding is counterintuitive. Void fraction is typically inversely correlated with density (see Supplementary Information Section 1.3). Although low density is generally associated with a low $\kappa$ value, as shown in Figure 2a, the unexpected observation is that a high void fraction (which corresponds to low density) surprisingly results in a high $\kappa$ value, as also depicted in Figure 2c.
Surface area: Gravimetric surface area (GSA), i.e., surface area per unit mass varies highly across all the COF structures (see Figure 2d) with a larger cluster of points in the 7000-8000 $\mathrm{m^{2}\ g^{-1}}$ range. We also observe that around intermediate GSA of 7,000 $\mathrm{m^{2}\ g^{-1}}$, we have the widest spread and highest values of $\kappa$.
Poor $\kappa$-prediction with known descriptors: From the above initial analysis, we notice that the geometrical descriptors used here are insufficient to provide solid trends with respect to thermal conductivity. For example, with all the optimal descriptor values, i.e., a large density, a low pore size, a large void fraction, and an intermediate GSA, we are still not guaranteed to obtain a COF with a high $\kappa$. To elucidate this further, we choose four COFs – denoted by (i)-(iv) – and compare them in Figure 2a-d.
COF (i) has the highest thermal conductivity in the dataset with $\kappa=4.025\ \mathrm{Wm^{-1}K^{-1}}$. The structure has a relatively small pore size with 13.71 Å (see Figure 2a), a medium density with 0.928 $\mathrm{g\ cm^{-3}}$ (see Figure 2b), a relatively high void fraction of 0.844 (see Figure 2c), and an intermediate surface area of 7381 $\mathrm{m^{2}g^{-1}}$. In contrast, the second COF (ii) possesses a similar pore size of 9.88 Å, a similar density of 0.938 $\mathrm{g\ cm^{-3}}$, a similar void fraction of 0.869, and a similar surface area of 7171 $\mathrm{m^{2}g^{-1}}$, but exhibits thermal transfer capabilities almost seven times lower than that of (i) with $\kappa=0.603\ \mathrm{Wm^{-1}K^{-1}}$.
The third COF (iii) has a pore size of 33.73 Å in the intermediate-to-high range, a low density at 0.408 $\mathrm{g\ cm^{-3}}$, a high void fraction at 0.932, and a slightly larger surface area of 8741 $\mathrm{m^{2}g^{-1}}$. Following the trends described in literature,[33, 42, 43] the geometrical descriptors are not optimal for a high $\kappa$. Nonetheless, COF (iii) has rather high $\kappa=1.96\ \mathrm{Wm^{-1}K^{-1}}$. Lastly, COF (iv) has an exceptionally high pore size at 80.68 Å and a low density of 0.262 $\mathrm{g\ cm^{-3}}$ but high void fraction of 0.958. Due to the extreme geometry, the COF (iv) has a very low thermal conductivity of $\kappa=0.289\ \mathrm{Wm^{-1}K^{-1}}$.
We further train four ensemble regression models as baselines based on Random Forest,[44] Gradient Boosting,[45] XGBoost,[46] and AdaBoost,[47] (see Supplementary Information Section 2 for training protocols and hyperparameters of each model) and features shown in Figure 3a to predict thermal conductivity. The same figure describes the corresponding goodness-of-fit $R^{2}$ for each model. The prediction accuracy is poor for all the models ($R^{2}<0.6$ across all models), and it is evident that these well-known descriptors are insufficient to predict the thermal conductivity. In summary, these examples illustrate the fact that the thermal conductivity structure-property relationships are highly complex and call for further examination.
Predicting thermal conductivity using deep learning
Figure 3: (a) Performance of ensemble regression models at predicting $\kappa$ using the shown descriptors. Average and standard goodness-of-fit $R^{2}$ are computed based on 10-fold cross-validation. (b) Schematic of PMTransformer model. A sample COF structure is shown on the left. A crystal graph convolutional neural network computes local embeddings of the COF graph, while the GRIDAY computes a three-dimensional energy grid of the structure, which becomes the global embeddings. These embeddings are combined and input into the PMTransformer (which includes a transformer encoder with attention mechanism and a prediction head) to predict $\kappa$. On the right, a parity plot compares $\kappa$ predictions with ground truth values. The dashed line represents the ideal line with zero intercept and unit slope.
Our starting point is the deep learning-based Porous Material Transformer (PMTransformer) developed by Park et al.[40] The PMTransformer is a multi-modal transformer[48] model (see Figure 3b for schematic) designed for universal transfer learning for predicting the properties of porous materials, including COFs, MOFs, and zeolites. It receives two distinct sets of inputs: local and global features. Local features, which reflect the chemistry of building blocks and bonds, are derived from a crystal graph convolutional neural network[49] (CGCNN). The CGCNN directly processes the crystal graph, where atoms and bonds are nodes and edges, respectively, and generates embeddings that capture essential chemical details. These embeddings serve as local features in the PMTransformer. Conversely, global features describe crystalline characteristics, including topological and geometric descriptors such as pore size and surface area. These features are derived from three-dimensional energy grids, which are generated by calculating the interaction energy between the material structure and a methane gas molecule at each grid point using the package GRIDAY[50]. Similar to images in Vision Transformers[51], the energy grids are divided into patches and flattened through linear projections. Eventually, both local and global embeddings are input into a transformer encoder, where the attention mechanism[48] helps the model focus on the most relevant parts of the molecular structure for predicting (in our case) the thermal conductivity.
To enable universal transfer learning, the transformer encoder in PMTransformer is pre-trained on an extremely large dataset of 1.9 million hypothetical porous materials to predict easily obtainable yet essential properties such as topology, void fraction, and building block prediction. This approach ensures the encoder captures critical information necessary for accurately predicting other, more complex properties in downstream tasks with much smaller datasets. For more details regarding the pre-training of PMTransfomer, refer to ref.[40]
In this study, we leveraged the pre-trained transformer encoder of the PMTransformer model to perform transfer learning for the prediction of the thermal conductivity of COFs. We fine-tune the transformer encoder model using the pre-trained weights as a starting point and jointly train a prediction head based on a multi-layer perceptron architecture to map the output of the transformer encoder to the thermal conductivity of COFs. We used the mean squared error (MSE) as a loss function for training. The dataset of 2,471 COFs is randomly split into two subsets: 90% of the data is used for training (out of which 10% for validation), and the remaining data is used for testing. Additional details of the training protocol are provided in Supplementary Information Section 3.
We demonstrate the model’s prediction accuracy on the test dataset in Figure 3b. The model achieves a goodness-of-fit $R^{2}$ of $0.909\pm 0.006$ (across five random seeds) in predicting $\kappa$, with a mean absolute error (MAE) of $0.075\ \mathrm{Wm^{-1}K^{-1}}$. Compared to the four baseline regression models, which rely on only hand-crafted and/or commonly used features (as shown in Figure 3a), the deep learning model significantly outperforms them all. The accurate predictions on the test dataset not only highlight the model’s ability to capture the underlying structure-property relationships governing the thermal conductivity of COFs but also hints at the potential for uncovering the existing knowledge gap in those structure-property relationships.
Figure 4: Two example pairs (contained in dashed boxes) of COFs with same topologies, similar geometric descriptors, but contrasting thermal conductivities. The first column illustrates the COF structures. The second column shows the atom-level attention score profile computed by the attention mechanism. The third column shows the same COF structure distinguishing atoms on the main branch and dangling atoms (with separate distinction for hydrogen atoms). The fourth column shows the VDOS profiles of various groups of atoms within the corresponding COF structure with the overlap metric $S$. The legend indicates the VDOS profile for main branch atoms (.)(.)( . ) and dangling atoms (.$(.)^{\text{(d)}}$.
Discovering novel thermal transfer mechanisms via self-attention
We leverage the multi-head self-attention mechanism[48] of the transformer architecture to interpret the predictions made by the deep learning model. The attention mechanism in transformers is designed to identify and assign weights, known as attention scores, to the significance of different parts of the input data/encodings relative to all the other parts. When applied to molecular structures, such as COFs, the attention mechanism enables the model to dynamically focus attention on key molecular substructures in relation to all the other substructures present based on their relevance to the property being predicted, in this case, thermal conductivity. In multi-head attention, multiple attention heads are needed to capture different aspects or relationships within the data simultaneously. Specifically, in the context of the PMTransformer model applied to a COF to predict its thermal conductivity, we calculate the attention for each atom in the COF by averaging the attention scores from all attention heads and computing the joint attention via multiplicative aggregation across all layers of the transformer encoder.
Additionally, we identify the main branch of a COF as the shortest continuous path connecting the boundary points necessary for periodicity. We then classify all the atoms extending and excluded from the main branch as dangling mass. More details on how we classify each atom as either part of the main branch or part of a dangling side branch are presented in Supplementary Information Section 4. We observe distinct patterns in the attention assigned to various atomic sites within a COF’s graph representation, which correspond to the location of the dangling masses. Two representative examples are shown in Figure 4.
Consider the first example (see Figure 4a-b). The top row illustrates a COF structure (Figure 4a) from the test dataset with a thermal conductivity of $1.105\ \mathrm{Wm^{-1}K^{-1}}$ and no dangling masses except hydrogen atoms. Additionally, the atom-wise attention scores are uniformly distributed and low for all the carbon atoms with some elevated attention to the nitrogen atoms. In the second row, we illustrate a COF structure (Figure 4b) with the same topology as the previous one, but with a much lower thermal conductivity of $0.533\ \mathrm{Wm^{-1}K^{-1}}$. Its attention profile reveals that certain atom groups are being paid special attention to by the transformer, i.e., specifically the $-\text{NO}_{2}$ functional group on the benzene ring that is dangling from the main branch of the COF. Notably, the same groups of atoms that exhibit higher attention scores are also classified as dangling masses (see Figure 4b attention profile & dangling mass).
[…]
The same trend is observed for numerous examples, with additional ones presented in the Supplementary Information Section 5. We conclude that this increased attention is indicative of the deep learning model’s understanding that these dangling masses are significant in predicting thermal conductivity. Furthermore, it suggests that the presence of dangling masses disrupts the heat transfer pathways and thereby reduces the thermal conductivity through the material. A similar effect of dangling mass lowering the thermal conductivity has been reported previously in singular polymer chains [52] and amorphous polymers, [53] but remains a novel predictor for COFs which are ordered polymeric structures.
To further validate the structure-property relationship interpreted from the deep learning model, we examine the impact of dangling atoms on the vibrational density of states (VDOS) for the aforementioned contrasting COF examples. VDOS, which characterizes the distribution of vibrational modes in a system as a function of frequency, is known to impact thermal transfer properties. Overlaps in VDOS profiles between different atoms have been shown to affect these properties in both COFs [54] and MOFs. [55]
Regression model: Random Forest Gradient Boosting XGBoost AdaBoost
Without DMR as feature $R^{2}$ $\mathbf{0.592\pm 0.039}$ $0.536\pm 0.038$ $0.551\pm 0.060$ $0.312\pm 0.103$
$R^{2}$ $\mathbf{0.728\pm 0.069}$ $0.688\pm 0.070$ $0.691\pm 0.035$ $0.416\pm 0.125$
Gini: Density $0.397$ $0.433$ $0.425$ $0.478$
Gini: DMR $0.278$ $0.286$ $0.296$ $0.243$
With DMR as feature Gini: LPD $0.185$ $0.218$ $0.146$ $0.084$
Gini: Void fraction $0.070$ $0.032$ $0.076$ $0.017$
Gini: Surface area $0.069$ $0.031$ $0.177$ $0.057$
Table 1: Comparison of performances of standard ensemble regression models at predicting $\kappa$ by excluding (top table) and including (bottom table) DMR as an input feature. We performed a 10-fold cross-validation to obtain the mean and standard deviations in performances. Gini Importance for all ensemble regression models (bottom table) for all features (including DMR) are listed below in descending order.
Through MD simulations, the VDOS is calculated using the Fourier transform of the normalized velocity autocorrelation function of specific groups of atoms (see Supplementary Information Section 6 for details). Figure 4 (third column) show the VDOS profile of carbon, nitrogen, and oxygen atoms in the representative COF examples. For any COF, we define a VDOS overlap metric $S$ as the ratio of the area under the curve (AUC) of the minimum VDOS across all atom types (at each frequency) and the AUC of the maximum VDOS across all atom types (at each frequency), i.e.,
$S=\frac{\int_{\omega}\min\left\{\ \mathrm{f^{V}_{C}(\omega),\ f^{V}_{N}(\omega),\ f^{V}_{O}(\omega),\ f^{V}_{B}(\omega),\ f^{V,d}_{C}(\omega),\ f^{V,d}_{N}(\omega),\ f^{V,d}_{O}(\omega),\ f^{V,d}_{B}(\omega)\ }\right\}\;\!\mathrm{d}\omega}{\int_{\omega}\max\left\{\ \mathrm{f^{V}_{C}(\omega),\ f^{V}_{N}(\omega),\ f^{V}_{O}(\omega),\ f^{V}_{B}(\omega),\ f^{V,d}_{C}(\omega),\ f^{V,d}_{N}(\omega),\ f^{V,d}_{O}(\omega),\ f^{V,d}_{B}(\omega)}\right\}\;\!\mathrm{d}\omega}.$ (1)
Here, $\omega$ denotes the frequency; $\mathrm{f_{C}^{V}}(\omega)$, $\mathrm{f_{N}^{V}}(\omega)$, $\mathrm{f_{O}^{V}}(\omega)$, and $\mathrm{f_{B}^{V}}(\omega)$ denote the VDOS at frequency $\omega$ for the carbon, nitrogen, oxygen and boron atoms, respectively. Additionally, we differentiate between atoms in the main branch and dangling atoms, where the VDOS profiles for dangling atoms are denoted with a (.$(.)^{\text{d}}$ superscript. A large overlap in VDOS profiles of different atom types would result in $S\approx 1$, whereas a small overlap would yield $S\approx 0$. A high overlap indicates that phonon waves (i.e., quantized modes of vibrations responsible for thermal energy transfer within the crystal lattice) can propagate freely throughout the structure. This would lead to minimized phonon scattering and facilitate efficient heat transfer across the material.[56]
In the COF structures, which lack or have minimal dangling atoms, the VDOS profiles of the atoms are broad and overlap significantly, e.g., $S\approx 0.41$ and $0.49$ for examples in Figure 4a and c, respectively. This results in an even distribution of vibrational modes across a wide frequency spectrum for all atoms. In contrast, the COFs with substantial dangling atoms exhibit a VDOS profile with minimal overlap, e.g., $S\approx 0.08$ and $0.09$ for examples in Figure 4b and d, respectively. The vibrational modes of individual atoms in these COFs are confined to much narrower frequency bands, which leads to a less harmonized VDOS profile. The lack of overlap suggests that the vibrations are highly localized around the dangling atoms, which causes the phonon waves to be scattered and disrupting their ability to transfer heat efficiently through the material.[57] We interpret the mismatch in the VDOS profiles as an energy barrier for phonon transport, where phonons encounter resistance, preventing smooth energy transfer across atoms. This suggests that the absence of dangling mass is crucial for enhancing thermal conductivity in COFs. Dangling atoms create these mismatches in VDOS, which in turn hinders phonon transport and reduces thermal conductivity.
To quantify the role of dangling masses in the structure-property relationship for thermal conductivity of COFs, we introduce the dangling mass ratio (DMR). The DMR $\in[0,1]$ quantifies the ratio between the mass of dangling atoms and the total mass of the COF, i.e.,
$\text{DMR}=\frac{\sum_{i\in\mathcal{D}}m_{{i}}}{\sum_{i\in\mathcal{D}}m_{{i}}+\sum_{i\in\mathcal{M}}m_{{i}}},$ (2)
where $m_{i}$ is the mass of the $i^{\text{th}}$ atom belonging to the set of either dangling atoms or of non-dangling main branch atoms – denoted by $\mathcal{D}$ and $\mathcal{M}$, respectively. Higher DMR values indicate a greater proportion of dangling masses relative to the COF’s total mass.
We then introduce DMR as an additional feature, alongside pore size, density, void fraction, and surface area, into the classical ensemble regression models that served as the baseline for PMTransformer. The inclusion of DMR results in a marked improvement in the goodness-of-fit (see Table 1) across all the models, i.e., random forest, gradient boosting, XGBoost and AdaBoost. To assess the importance of individual features, including DMR, we analyzed their Gini importance.[58] These values quantify the contribution of each feature to the overall predictive power of the model by measuring how much each feature reduces the Gini impurity or randomness when making decisions in the ensemble model. We note that DMR emerges as the second most influential predictor with 24-30% importance, closely following behind density (see Table 1). This underscores the contribution of the DMR to the models’ predictive capabilities and the role of dangling atoms as a novel predictor for thermal conductivity in COFs.
Generalizability and high-throughput screening for discovering COFs with tailored properties
Figure 5: High-throughput screening. (a) Distribution of $\kappa$ in the training dataset. Dashed lines indicated minimum, mean, and maximum $\kappa$ across the training dataset. (b,c) Distribution of $\kappa$ of selected COF candidates identified in high-throughput screening for low and high $\kappa$ in the unlabeled datasets of (b) Mercado et al.[41] and (c) Lan et al.[59] While the former dataset shares the same design space as the training dataset, the latter has a different design space. All thermal conductivities shown here are computed using MD. (d) Representative COFs identified through high-throughput screening with $\kappa$ lower and higher than the minimum and maximum, respectively, across the training dataset. Also indicated are their corresponding DMR and $S$ values.
To evaluate the generalization capabilities of the deep learning model for high-throughput screening, we screen previously unseen 6,170 two-dimensional COFs in the unlabelled dataset by Mercado et al.[41] from which our training and test sets were originally subsampled. Moreover, we screen $35,638$ two-dimensional COFs in the unlabelled dataset by Lan et al.[59] This dataset covers a vastly different design space with a variety of linkers and knots previously unseen by our trained model.
The screening is performed with the objective of facilitating the identification COFs with extremely high or low thermal conductivity without resorting to computationally expensive simulations. We observe an average screening rate of 0.07 seconds per COF structure (see Supplementary Information Section 7 for details on computational runtimes), thereby enabling a speedup of almost seven orders of magnitude compared to MD-based screening.
Once potential candidates with extreme thermal conductivities are identified using the PMTransformer, we then calculate their thermal conductivities through MD simulations for these selected few, rather than for the entire datasets. The results are summarized in Figure 5. Our screening process on the dataset by Mercado et al.[41] revealed COFs exhibiting thermal conductivities up to twice the maximum value observed in the training set, but stemming from the same design space. By screening the dataset by Lan et al.,[59] we discover COFs with thermal conductivities five times as high as the ones in our training set. Additionally, we discovered several COFs with thermal conductivities lower than any values previously recorded in our dataset. A visual inspection reveals that these COFs feature a combination of large pore sizes, low density, and a substantial presence of dangling masses. For all the identified COFs, we observe that thermal conductivity shows a strong trend with both dangling mass ratio (DMR) and VDOS overlap ratio ($S$), which is in agreement with the previous observations.
Conclusion
In this study, we compiled an extensive dataset of thermal conductivities for 2,471 two-dimensional COFs, computed using NEMD. Despite observing some structure-property trends with conventional descriptors such as density, pore size, void fraction, and surface area, no single descriptor or combination thereof consistently predicted thermal conductivity with high accuracy, highlighting the complexity of COF structure-property relationships. To enhance prediction accuracy, we trained a transformer-based deep learning model incorporating a multi-head attention mechanism. This model significantly outperformed traditional ensemble regression models, achieving an $\mathrm{R^{2}}$ of $0.908\pm 0.007$.
Further analysis using the transformer’s attention mechanism revealed that COFs with higher amounts of dangling atoms exhibited lower thermal conductivities due to disrupted heat transfer pathways, a novel and significant predictor of thermal conductivity not previously reported. Additionally, incorporating the dangling mass ratio, i.e., DMR as an extra descriptor in traditional ensemble regression models significantly improved prediction accuracy. Across different models, DMR exhibited 24-30% in Gini importance, underscoring the importance of dangling mass in structure-property relationships. The vibrational density of states analysis corroborated these findings, showing that dangling masses create mismatched vibrational modes that hinder thermal transfer.
We further leveraged the deep learning model to efficiently screen thousands of COFs, both within and beyond the design space we analyzed, to identify candidates with extreme thermal conductivities. This approach provides a rapid and reliable method as well as valuable insights for designing COFs with tailored thermal characteristics. To support further research, we release the dataset of COF thermal conductivities and encourage the research community to utilize and expand upon it. Finally, we emphasize the need to bridge high interpretability with high accuracy in machine learning models to navigate the complex structure-property space of COFs and other nanoporous materials, thereby enabling the design of optimal materials for applications such as gas separation and thermal management.
Code availability
The codes generated during the current study are available at <link will be provided at time of publication>.
Data availability
The datasets generated during the current study are available at <link will be provided at time of publication>.
Author contributions
P.T. and Y.Z. generated the dataset. Y.Z. developed the NEMD pipeline to compute the thermal conductivities. Y.G. developed the dangling mass computation algorithm. P.T. and Y.Z. analyzed the data. P.T. trained the PMTransformer model and analyzed the results. P.T., Y.Z., and Y.G. visualized the results. S.K. and A.V. conceived and supervised the research. All authors contributed to writing the manuscript and approved the final version.
Competing interests
The authors declare no competing interests.
Acknowledgements
P.T., Y.G., and S.K. acknowledge that this material is based upon work supported by the Air Force Office of Scientific Research under award number FA8655-23-1-7020. P.T., Y.G., and S.K. acknowledge that this work was supported in part by Oracle Cloud credits and related resources provided by Oracle for Research. A.V. and Y.Z. acknowledge Microsoft Climate Research Initiative and University of Washington for funding. We thank Prof. Marcel Sluiter of TU Delft for his helpful discussions and insights during the course of this research. Additionally, we thank Eric Vo of University of Washington for his help in preparing simulations of COF structures.